# SCAWaveNet: A Spatial-Channel Attention-based Network for Global Significant Wave Height Retrieval

### v1

@@ Abstract
Recent advancements in spaceborne GNSS missions have produced extensive global datasets, providing a robust basis for deep learning-based significant wave height (SWH) retrieval. While existing deep learning models predominantly utilize CYGNSS data with four-channel information, they often adopt single-channel inputs or simple channel concatenation without leveraging the benefits of cross-channel information interaction during training. To address this limitation, a novel spatial–channel attention-based network, namely SCAWaveNet, is proposed for SWH retrieval. Specifically, features from each channel of the DDMs are modeled as independent attention heads, enabling the fusion of spatial and channel-wise information. For auxiliary parameters, a lightweight attention mechanism is designed to assign weights along the spatial and channel dimensions. The final feature integrates both spatial and channel-level characteristics. Model performance is evaluated using four-channel CYGNSS data. When ERA5 is used as a reference, SCAWaveNet achieves an average RMSE of 0.438 m. When using buoy data from NDBC, the average RMSE reaches 0.432 m. Compared to state-of-the-art models, SCAWaveNet reduces the average RMSE by at least 3.52% on the ERA5 dataset and by 5.47% on the NDBC buoy observations. The code is available at https://github.com/Clifx9908/SCAWaveNet.
@@ I Introduction
As a critical indicator of ocean wave energy and sea state conditions, Significant Wave Height (SWH) plays a key role in marine engineering, maritime safety, and climate change research[1]. Accurate retrieval of SWH provides a fundamental reference for the design of offshore structures, effectively mitigating the risk of structural failure under extreme sea conditions[2]. Precise SWH data also optimizes maritime route planning, enhancing safety and fuel efficiency[3]. In addition, long-term SWH monitoring aids in identifying extreme events and assessing marine resources, supporting sustainable ocean development[4, 5].
In recent years, Spaceborne Global Navigation Satellite System Reflectometry (GNSS-R) has become a mainstream technology for SWH retrieval[6]. Key missions include TechDemoSat-1[7], Cyclone GNSS (CYGNSS)[8], BuFeng-1 A/B[9], Fengyun-3E/3F/3G[10]. Unlike traditional methods[11, 12, 13], GNSS-R provides global, high-resolution SWH observations without requiring costly in situ instruments, significantly reducing operational costs. It remains robust under extreme sea conditions, providing reliable data in various weather and illumination conditions[14, 15, 16, 17]. This technique improves the accuracy of the SWH retrieval and broadens the scope of scientific and operational marine applications.
@@
As a powerful tool in GNSS-R observation, machine learning (ML) techniques have been widely applied in SWH retrieval[18, 19, 20]. Traditional ML approaches, such as Bagging Tree (BT)[21], Random Forest (RF)[22] and eXtreme Gradient Boosting (XGBoost)[23], typically construct regression models based on auxiliary parameters (APs). While these models demonstrate promising performance, they neglect the valuable information in two-dimensional Delay-Doppler maps (DDMs), thereby constraining SWH retrieval accuracy.
@@
Recent advancements in deep learning (DL) models have enabled high-resolution global SWH retrieval through architectures like Multi-scale Conv-BiLSTM[24], CNN-ConvLSTM-FCN[25], WaveTransNet[26], ViT-Wave[27]. In these methods, dual-branch architectures are widely adopted, where Convolutional Neural Networks (CNNs) extract features from DDMs in parallel with Multilayer Perceptrons (MLPs) processing APs. However, the fixed-weight nature of CNNs and MLPs makes it difficult to adjust features in various sea conditions, and they are limited in capturing long-range spatial dependencies. Inspired by Transformer[28] and Vision Transformer (ViT)[29], recent studies have incorporated attention mechanisms to enhance spatial feature flexibility and dependency modeling. Yet, most methods either utilize single-channel CYGNSS data or naively aggregate all channels, overlooking inherent channel-wise information and restricting architectural optimization potential.
@@
CYGNSS data inherently contain four-channel information. As depicted in Fig. 1(a), each receiver satellite in the CYGNSS constellation carries a four-channel Delay-Doppler Mapping Instrument (DDMI), capable of simultaneously capturing direct GPS signals and four reflected signals[30, 31, 32]. The direct signal enables precise receiver localization, and the reflected signals capture sea surface roughness for SWH estimation. During signal propagation, the specular point locations of each channel are spatially close, leveraging this spatial similarity can improve model’s predictive performance[33]. In addition, CYGNSS datasets include diverse DDMs (e.g., effective scattering area) and APs (e.g., specular point incidence angle, leading edge slope). Notably, the four channels exhibit similarity in DDM/AP characteristics due to their aligned spatial coverage and synchronous acquisition (Fig. 1(b)-(d)).
@@
To further quantify the intra-channel similarity, we analyzed the SWH density distribution from CYGNSS-ERA5 data in 2022, along with the standard deviation (SD) distribution across channels. As shown in Fig. 1(e), the SWH distributions across the channels are highly consistent, with most data concentrated between 1-3 m. Furthermore, Fig. 1(f) reveals that the 95th percentile of the SWH SD is 0.565 m, indicating minimal inter-channel variability for most samples. These findings confirm the similarity of reference values across CYGNSS channels, motivating the development of channel-aware models.
@@
Inspired by CYGNSS’s multi-channel observation mechanism, we propose SCAWaveNet, a spatial–channel attention (SCA) based model for SWH retrieval. Unlike prior works, SCAWaveNet introduces a novel four-channel modeling paradigm by jointly processing DDMs and APs from all channels. First, an SCA-based Transformer Encoder is employed to extract DDM features. Specifically, the DDMs are embedded, and the embeddings from each channel are treated as independent attention heads in multi-head self-attention block, enabling cross-dimensional information fusion. In parallel, a lightweight dual-dimensional attention mechanism is applied to generate spatial-channel weights for APs, and produce channel-aware features. Finally, these two types of features are fused and processed through a task-specific head to generate SWH estimates for all channels.
In summary, the contributions of this work are as follows.
@@
  • We propose a data processing framework for four-channel CYGNSS data and perform a comprehensive evaluation of four-channel information utilization for SWH retrieval. Experiments on the ERA5 and NDBC buoy datasets validate the effectiveness and robustness of SCAWaveNet.
  • We analyze architectural differences between single-channel and four-channel models, examining the impact of channel independence/dependence strategies on retrieval accuracy. Additional ablation studies clarify the role of key design choices in SCAWaveNet.
@@ II Related Work
In this section, we present a brief review of machine learning-based SWH retrieval models and introduce channel strategies in deep learning, emphasizing the importance of applying a spatial-channel attention mechanism in SWH retrieval.
@@ II-A Machine Learning Approaches for SWH Retrieval Task
Based on the Transformer and ViT architectures, researchers proposed WaveTransNet[26] and ViT-Wave [27], incorporating attention mechanisms. Both apply Transformer encoders to DDM patch embeddings but differ in AP branch design: ViT-Wave directly concatenates DDM and AP features, while WaveTransNet uses attention-based AP feature extraction before fusion. These attention mechanisms effectively capture long-range dependencies and enhance dynamic feature extraction. However, these models remain “single-channel” in paradigm, relying on either single-channel data or merged multi-channel inputs to predict a single SWH value. This approach neglects channel-wise information inherent in multi-channel GNSS-R data. Channel-aware modeling could optimize architectures, improve retrieval accuracy, and expand applications.
@@ II-B Channel Strategies in Deep Learning
Recently, channel strategies have emerged as a key research direction in multivariate prediction. Leveraging channel correlations can significantly enhance prediction accuracy, particularly for specific channels. For instance, in precipitation forecasting, by integrating multi-source data (e.g., temperature, humidity, atmospheric pressure, and wind speed), meteorological models can more accurately identify the occurrence and intensity of rainfall. The interconnections among different factors provide a more comprehensive framework for modeling rainfall dynamics [36].
Figure 2: Overview of channel strategies.
The concept of channel strategies has been introduced in multivariate time series forecasting[37, 38]. As illustrated in Fig. 2, existing channel strategies can be classified into three types:
@@
These strategies are often embedded in DL networks to enhance model performance. For example, SENet[39] pioneered the channel attention mechanism via the Squeeze-and-Excitation block, effectively modeling channel-wise dependencies to boost the representational ability of CNNs. Similarly, iTransformer[40] integrates channel information into the Transformer encoder, treating individual time series as tokens and applying self-attention mechanisms to capture inter-variable relationships, thereby improving the generalization ability of Transformers across different variables.
SWH retrieval is also a multivariate prediction task, which essentially examines the impact of SWH-related variables on SWH prediction. While existing deep learning-based SWH models primarily utilize CYGNSS data, they often overlook inherent channel correlations despite high inter-channel similarity. This paper proposes, for the first time, a four-channel modeling paradigm in SWH retrieval. We compare the characteristics of single-channel and four-channel models, and explore the effects of CI and CD strategies on the performance of the four-channel model.
@@ III Methodology
This section provides a comprehensive overview of the SCAWaveNet architecture, as illustrated in Fig. 3. The model consists of three core components: (1) DDM Branch: Utilizes a Transformer Encoder to apply spatial and channel attention mechanisms sequentially to four-channel DDM embeddings; (2) APs Branch: Employs a lightweight attention module to compute cross-dimensional attention weights and integrate these into auxiliary parameters; (3) Feature Fusion and Task Head: Fuses multi-branch features to generate four-channel SWH predictions. Each component will be detailed in the following subsections.
@@ III-A Transformer-based DDM branch
This branch is designed to effectively capture the spatial and channel-wise features of DDMs. Previous studies have not sufficiently leveraged the intrinsic features of DDMs. On the one hand, existing approaches typically adopt CNNs as feature extraction modules, which is limited in feature representation ability. On the other hand, these approaches focused only on spatial information while overlooking channel-wise information in DDMs. To address these challenges, we integrate the Spatial-Channel Attention (SCA) mechanism into the DDM branch, which includes three components: an embedding block, a channel aggregation block, and an SCA-based Transformer encoder. These components are elaborated in the subsequent sections.
@@ III-A1 Embedding
Given the DDMs of each channel $x\in\mathbb{R}^{T\times W\times H}$, where $T$ denotes the number of DDM types, i.e., BRCS DDM (Brcs), effective scatter area DDM (Scatter) and analog power DDM (Power). Some models treat different types of DDM as “channels”[41, 26, 42], but we treat the number of signal paths received simultaneously by the receiver as channels. $W$ and $H$ represent the frequency shift and time delay dimensions of the DDMs, respectively. Following the patch embedding strategy in ViT[29], the DDMs are divided into non-overlapping patches using a 2D convolutional layer. Automatic padding is applied to incomplete regions. This process is formulated as:
@@
where, $x_{glb}$ is treated as a learnable parameter dynamically updated during training.
@@
$pos$ is the position, $d$ denotes the index of the dimension, and $d_{model}$ is the representation dimension of Transformer. Each even-indexed dimension is encoded with a sine function, and each odd dimension is encoded with a cosine function. $PE(x_{p}^{\prime})$ denotes the positional encoding for DDMs and $E_{p}$ is the final patch embedding in each channel after incorporating positional information.
@@ III-A3 Transformer Encoder
The MSA module is the core component of the Transformer Encoder. Related studies did not consider channel information in DDM embedding, leading to limited representation of DDMs. To address this, we propose an enhanced MSA block that integrates the SCA mechanism to enable richer feature extraction from DDMs, as shown in Fig. 4. Specifically, the input embedding $E$ is divided into three equal vectors: query $Q\in\mathbb{R}^{M\times d_{\text{model}}}$, key $K\in\mathbb{R}^{M\times d_{\text{model}}}$, and value $V\in\mathbb{R}^{M\times d_{\text{model}}}$. They are projected by the learnable weight matrix $W^{Q}$, ${W}^{K}$, and $W^{V}$, resulting in transformed representations:
@@
where $m$ indexes the tokens in the sequence $M$, ${H}[m,i]$ represents the output of the $m$-th token from the $i$-th channel, and ${W}^{O}[i,j]$ controls the contribution of the $i$-th channel to the $j$-th output dimension. If the weights in ${W}^{O}$ are non-zero, channel attention is introduced, influencing every element of the output. Therefore, the degree of channel interaction is closely related to ${W}^{O}$. The projection matrix ${W}^{O}$ is optimized during backpropagation to balance channel interactions. To ensure accurate SWH prediction from all channels, its weights are updated via gradient descent to balance the contributions of each channel to the final output.
The Add & Norm component consists of two primary operations: residual connection (RC) and layer normalization (LN). For the output of the MSA, LN is first applied to normalize the features of each sample, stabilizing the training process. Then a dropout operation is introduced to mitigate overfitting. Finally, an RC is applied by adding the normalized features to the original MSA output. This procedure is formulated as:
@@
The output ${D}$ is then fed into an FFN, followed by another Add & Norm operation, to further enhance the feature representation. The FFN consists of two linear layers and a dropout operation. Specifically, ${D}$ is first projected through a linear layer, followed by ReLU activation and dropout function, and then mapped by a second linear layer to produce the FFN output ${D}_{f}$. Then ${D}_{f}$ is processed by another Add & Norm component with ${D}$, producing the output of the encoder layer:

### v2

@@ Abstract
Recent advancements in spaceborne GNSS missions have produced extensive global datasets, providing a robust basis for deep learning-based significant wave height (SWH) retrieval. While existing deep learning models predominantly utilize CYGNSS data with four-channel information, they often adopt single-channel inputs or simple channel concatenation without leveraging the benefits of cross-channel information interaction during training. To address this limitation, a novel spatial–channel attention-based network, namely SCAWaveNet, is proposed for SWH retrieval. Specifically, features from each channel of the DDMs are modeled as independent attention heads, enabling the fusion of spatial and channel-wise information. For auxiliary parameters, a lightweight attention mechanism is designed to assign weights along the spatial and channel dimensions. The final feature integrates both spatial and channel-level characteristics. Model performance is evaluated using four-channel CYGNSS data. When ERA5 is used as a reference, SCAWaveNet achieves an average RMSE of 0.438 m. When using buoy data from NDBC, the average RMSE reaches 0.432 m. Compared to state-of-the-art models, SCAWaveNet reduces the average RMSE by at least 3.52% on the ERA5 dataset and by 5.68% on the NDBC buoy observations. The code is available at https://github.com/Clifx9908/SCAWaveNet.
@@ I Introduction
As a critical indicator of ocean wave energy and sea state conditions, Significant Wave Height (SWH) plays a key role in marine engineering, maritime safety, and climate change research [1]. Accurate retrieval of SWH provides a fundamental reference for the design of offshore structures, effectively mitigating the risk of structural failure under extreme sea conditions [2]. Precise SWH data also optimizes maritime route planning, enhancing navigation safety and fuel efficiency [3]. In addition, long-term SWH monitoring aids in identifying extreme events and assessing marine resources, supporting sustainable ocean development [4, 5].
In recent years, Spaceborne Global Navigation Satellite System Reflectometry (GNSS-R) has become a mainstream technology for SWH retrieval [6]. Key missions include TechDemoSat-1 [7], Cyclone GNSS (CYGNSS) [8], SMAP-R [9], Fsscat [10], Spire [11], Fengyun-3E/3F/3G [12], BuFeng-1 A/B [13], and Tianmu-1 [14, 15]. Unlike traditional methods [16, 17, 18], GNSS-R enables global, high-resolution SWH observations without relying on costly in situ instruments, substantially reducing operational costs. It remains robust under extreme sea conditions, providing reliable data in various weather and illumination scenarios [19, 20, 21, 22]. This technique improves the accuracy of SWH retrieval and expands the scope of scientific and operational marine applications.
@@
As a powerful tool in GNSS-R observation, machine learning (ML) techniques have been widely applied to SWH retrieval [23, 24, 25]. Traditional ML approaches, such as Bagging Tree (BT) [26], Random Forest (RF) [27] and eXtreme Gradient Boosting (XGBoost) [28], typically construct regression models based on auxiliary parameters (APs). While these models demonstrate promising performance, they neglect the valuable information in two-dimensional Delay-Doppler maps (DDMs), thus constraining SWH retrieval accuracy.
@@
Recent advancements in deep learning (DL) models have enabled high-resolution global SWH retrieval through architectures like Multi-scale Conv-BiLSTM [29], CNN-ConvLSTM-FCN [30], WaveTransNet [31], and ViT-Wave [32]. In these methods, dual-branch architectures are widely adopted, where Convolutional Neural Networks (CNNs) extract features from DDMs in parallel with Multilayer Perceptrons (MLPs) processing APs. However, the fixed-weight nature of CNNs and MLPs hinders their ability to adjust features in various sea conditions, and they are limited in capturing long-range spatial dependencies. Inspired by Transformer [33] and Vision Transformer (ViT) [34], recent studies have incorporated attention mechanisms to enhance spatial feature flexibility and dependency modeling. Yet, most methods either rely on single-channel CYGNSS data or naively aggregate all channels, overlooking inherent channel-wise information and thus restricting architectural optimization potential.
@@
CYGNSS data inherently contain four-channel information. As depicted in Fig. 1(a), each receiver satellite in the CYGNSS constellation carries a four-channel Delay-Doppler Mapping Instrument (DDMI), capable of simultaneously capturing direct GPS signals and four reflected signals [35, 36, 37]. The direct signal enables precise receiver localization, the reflected signals capture sea surface roughness for SWH estimation. During signal propagation, the specular point locations of each channel are spatially close, leveraging this spatial similarity can improve the model’s predictive performance [38]. In addition, CYGNSS datasets include diverse DDMs (e.g., effective scattering area) and APs (e.g., specular point incidence angle, leading edge slope). Notably, the four channels exhibit similarity in DDM/AP characteristics due to their aligned spatial coverage and synchronous acquisition (Fig. 1(b)-(d)).
@@
To further quantify the intra-channel similarity, we analyzed the SWH density distribution derived from CYGNSS-ERA5 data in 2022, along with the standard deviation (SD) distribution across channels. As shown in Fig. 1(e), the SWH distributions are highly consistent among channels, with most data concentrated between 1-3 m. Furthermore, Fig. 1(f) reveals that the 95th percentile of the SWH SD is 0.565 m, indicating minimal inter-channel variability for the majority samples. These findings confirm the similarity of reference values across CYGNSS channels, motivating the development of channel-aware modeling approaches.
@@
Inspired by CYGNSS’s multi-channel observation mechanism, we propose SCAWaveNet, a Spatial–Channel Attention (SCA) based model for SWH retrieval. Unlike prior works, SCAWaveNet introduces a novel four-channel modeling paradigm by jointly processing DDMs and APs from all channels. First, an SCA-based Transformer Encoder is employed to extract DDM features. Specifically, the DDMs are embedded, and the embeddings from each channel are treated as independent attention heads in multi-head self-attention block, enabling cross-dimensional information fusion. In parallel, a lightweight dual-dimensional attention mechanism is applied to generate spatial-channel weights for APs, and produce channel-aware features. Finally, these two types of features are fused and passed through a task-specific head to generate SWH estimates for all channels.
In summary, the contributions of this work are as follows:
@@
  • We design a data processing framework for four-channel CYGNSS data and conduct a comprehensive evaluation of four-channel information utilization for SWH retrieval. Experiments on the ERA5 and NDBC buoy datasets validate the effectiveness and robustness of SCAWaveNet.
  • We analyze architectural differences between single-channel and four-channel models, investigating the impact of channel independence/dependence strategies on retrieval accuracy. Additional ablation studies further clarify the role of key design choices in SCAWaveNet.
@@ II Related Work
In this section, we present a brief review of machine learning-based SWH retrieval models and introduce channel strategies in deep learning, emphasizing the significance of employing a spatial-channel attention mechanism in SWH retrieval.
@@ II-A Machine Learning Approaches for SWH Retrieval Task
Based on the Transformer and ViT architectures, researchers proposed WaveTransNet [31] and ViT-Wave [32], incorporating attention mechanisms. Both apply Transformer encoders to DDM patch embeddings but differ in AP branch design: ViT-Wave directly concatenates raw APs with DDM features, while WaveTransNet uses attention-based AP feature extraction before two-branch fusion. These attention mechanisms effectively capture long-range dependencies and enhance dynamic feature modeling. However, these models remain “single-channel” in paradigm, relying on either single-channel data or merged multi-channel inputs to predict a single SWH value. This approach neglects channel-wise information inherent in multi-channel GNSS-R data. To resolve this issue, channel-aware modeling could optimize architectures, improve retrieval accuracy, and expand applications.
@@ II-B Channel Strategies in Deep Learning
Figure 2: Overview of existing channel strategies.
Recently, channel strategies have emerged as a key research direction in multivariate prediction. Leveraging channel correlations can significantly enhance prediction accuracy, especially for specific channels. For instance, in precipitation forecasting, by integrating multi-source data (e.g., temperature, humidity, atmospheric pressure, and wind speed), meteorological models can more accurately identify the occurrence and intensity of rainfall. The interconnections among different factors provide a more comprehensive framework for modeling rainfall dynamics [41].
The concept of channel strategies has been introduced in multivariate time series forecasting [42, 43]. As illustrated in Fig. 2, channel strategies can be classified into three types:
@@
These strategies are often embedded in DL networks to enhance model performance. For example, SENet [44] pioneered the channel attention mechanism via the Squeeze-and-Excitation block, effectively modeling channel-wise dependencies to boost the representation ability of CNNs. Similarly, iTransformer [45] integrates channel information into the Transformer encoder, treating individual time series as tokens and applying self-attention mechanisms to capture inter-variable relationships, thereby improving the generalization ability of Transformers across different variables.
SWH retrieval is also a multivariate prediction task, which essentially investigates the influence of SWH-related variables in SWH prediction. While existing deep learning-based SWH models primarily utilize CYGNSS data, they often overlook inherent channel correlations despite high inter-channel similarity. This paper proposes, for the first time, a four-channel modeling paradigm for SWH retrieval. We compare the characteristics of single-channel and four-channel models, and explore the effects of CI and CD strategies on the performance of the four-channel model.
@@ III Methodology
This section provides a comprehensive overview of the SCAWaveNet architecture, as illustrated in Fig. 3. The model consists of three core components: (1) DDM Branch: employs a Transformer Encoder to apply spatial and channel attention mechanisms sequentially to four-channel DDM embeddings; (2) AP Branch: utilizes a lightweight attention module to compute cross-dimensional attention weights and integrate them into auxiliary parameters; (3) Feature Fusion and Task Head: Fuses multi-branch features to generate four-channel SWH predictions. Each component will be detailed in the following subsections.
@@ III-A Transformer-based DDM branch
This branch is designed to effectively capture the spatial and channel-wise features of DDMs. Previous studies have not sufficiently leveraged the intrinsic features of DDMs. On the one hand, existing approaches typically adopt CNNs as feature extraction modules, which are limited in feature representation ability. On the other hand, these approaches focus only on spatial information while overlooking channel-wise information in DDMs. To address these challenges, we integrate the Spatial-Channel Attention (SCA) mechanism into the DDM branch, which includes three components: an embedding block, a channel aggregation block, and an SCA-based Transformer encoder. These components are elaborated in the subsequent sections.
@@ III-A1 Embedding
Given the DDMs of each channel $x\in\mathbb{R}^{T\times W\times H}$, where $T$ denotes the number of DDM types, i.e., BRCS DDM (Brcs), effective scatter area DDM (Scatter) and analog power DDM (Power). Some models treat different types of DDM as “channels”[46, 31, 47], but we treat channels as the number of signal paths simultaneously received by the receiver. $W$ and $H$ represent the frequency shift and time delay dimensions of the DDMs, respectively. Following the patch embedding strategy in ViT [34], the DDMs are divided into non-overlapping patches using a 2D convolutional layer. Automatic padding is applied to handle incomplete regions. This process is formulated as:
@@
where, $x_{glb}$ is defined as a learnable parameter dynamically updated during training.
@@
$pos$ is the position, $d$ is the dimension index, $d_{model}$ is the representation dimension of Transformer. Each even-indexed dimension is encoded with a sine function, while each odd dimension is encoded with a cosine function. $PE(x_{p}^{\prime})$ denotes the positional encoding for DDMs. $E_{p}$ denotes the final patch embedding in each channel after incorporating positional information.
@@ III-A3 Transformer Encoder
The MSA module is the core component of the Transformer Encoder. Related studies failed to consider channel information in DDM embedding, leading to limited representation of DDMs. To address this, we propose an enhanced MSA block that integrates the SCA mechanism to enable richer feature extraction from DDMs, as shown in Fig. 4. Specifically, the input embedding $E$ is divided into three equal vectors: query $Q\in\mathbb{R}^{M\times d_{\text{model}}}$, key $K\in\mathbb{R}^{M\times d_{\text{model}}}$, and value $V\in\mathbb{R}^{M\times d_{\text{model}}}$. They are projected by the learnable weight matrix $W^{Q}$, ${W}^{K}$, and $W^{V}$, resulting in transformed representations:
@@
where $m$ indexes the tokens in the sequence $M$, ${H}[m,i]$ represents the output of the $m$-th token from the $i$-th channel, and ${W}^{O}[i,j]$ determines the contribution of the $i$-th channel to the $j$-th output dimension. If the weights in ${W}^{O}$ are non-zero, channel attention is introduced, which affects every element of the output. Therefore, the degree of channel interaction is closely related to ${W}^{O}$. The projection matrix ${W}^{O}$ is optimized during backpropagation to dynamically adjust channel interactions. To ensure accurate SWH prediction from all channels, its weights are updated via gradient descent to balance the contributions of each channel to the final output.
The Add & Norm component consists of two primary operations: residual connection (RC) and layer normalization (LN). For the output of the MSA, LN is first applied to normalize the features of each sample, stabilizing the training process. Then a dropout operation is introduced to mitigate overfitting. Finally, an RC is applied by adding the normalized features back to the original MSA output. This procedure is formulated as:
@@
The output ${D}$ is then fed into an FFN, followed by another Add & Norm operation, to further refine the feature representation. The FFN consists of two linear layers and a dropout operation. Specifically, ${D}$ is first projected through a linear layer, followed by ReLU activation and a dropout function, and then mapped by a second linear layer to produce the FFN output ${D}_{f}$. Then, ${D}_{f}$ is processed by another Add & Norm component with ${D}$, producing the output of the encoder layer: